OpenStudy (jennyrlz):

help!!!!! Some college discrete math :3

OpenStudy (jennyrlz):

From the given functions defined from R to R, identify the bijective functions. f(x) = –3x + 4

OpenStudy (jennyrlz):

give me a sec, there are more choice ^.^

OpenStudy (mathmate):

This question requires an understanding of the meaning of the word "bijective". Once that's understood, the answer will be obvious. Bijection (according to Wiki) is "In mathematics, a bijection, bijective function or one-to-one correspondence is a function between the elements of two sets, where every element of one set is paired with exactly one element of the other set, and every element of the other set is paired with exactly one element of the first set." If you google Bijective function, the heading gives a very nice visual illustration of exactly what bijection means.

OpenStudy (jennyrlz):

Fx=-3x^2+7 Fx=(x+1)/(x+2) X^5+1=fx

OpenStudy (jennyrlz):

Ok thanks a lot I'll look into bike give xD

OpenStudy (jennyrlz):

Bijective*

OpenStudy (jennyrlz):

I'm a little lost so I apreciate the help :)

OpenStudy (mathmate):

This is the description I was talking about. http://prntscr.com/8o4ozw In short, a bijective function has to satisfy both the vertical line test and horizontal line test, which you shoud have learned about in functions. From you understanding of the functions, you can sketch each and find out which one satisfies both the hor. and vert. line test. (also bear in mind that ALL functions should satisfy the vertical line test, so that cuts your work by two.)

OpenStudy (jennyrlz):

I found out thanks to you first reply :). Now I actually understand what I am doing ^.^

OpenStudy (jennyrlz):

Your*

OpenStudy (jennyrlz):

If I have anymore questions I will ask you ^.^

OpenStudy (mathmate):

Very well! Any time!